For a species with a diploid number of 26, indicate how many chromosomes will be present in the somatic nuclei of individuals that are haploid, triploid, tetraploid, trisomic, and monosomic.

Part A: Haploid

Express your answer as an integer.

Part B: Triploid

Express your answer as an integer.

Part C: Tetraploid

Express your answer as an integer.

Part D: Trisomic

Express your answer as an integer.

Part E: Monosomic

Express your answer as an integer.
Answer #1

Answer:-

The term 'ploidy' refers to the number of sets of chromosomes found in the nucleus of the cell. The number of chromosomes in a set is represented by 'x'.

If the nucleus of the cell consists only set of chromosomes, then it is called haploid (x). If contains two sets of chromosomes, then it is diploid (2x). If three sets, then it is a triploid (3x).

Part A:-

Diploid (2x) = 26

therefore, x = 26/2 = 13

Thus, haploid (x) contains 13 chromosomes.

Part B:-

Triploid = 3x = 3 × 13 = 39 chromosomes.

​​​​​​Part C:-

Tetraploid = 4x = 4 × 13 = 52 chromosomes.

Part D:-

Trisomy are those diploids which contains an extra chromosome = (2x + 1) = (26 + 1) = 27 chromosomes.

Part E:-

Monosomy = (2x -1) = (26 - 1) = 25 chromosomes.

--------------------------------------------------